Could someone please help me with this problem?
Jo Ellen found that if the product of a number and 5 is increased by 20, the result is 28 less than the opposite of the number. What is the number?

Answer by ankor@dixie-net.com(22740)

Jo Ellen found that if the product of a number and 5 is increased by 20, the result is 28 less than the opposite of the number.
What is the number?
:
let x = "the number"
then
-x = "the opposite of the number"
:
Just write an equation for what it says
5x + 20 = -x - 28
5x + x = -28 - 20
6x = -48
x = -48/6
x = -8, it's opposite would be +8
;
:
Check this solution in the original equation
5(-8) + 20 = +8 - 28
-40 + 20 = -20
-20 = -20